OpenStudy (anonymous):

cube root of -512/343 -

OpenStudy (theeric):

Are you allowed to use a calculator?

OpenStudy (theeric):

He typed "cube root," @mayankdevnani ! :)

OpenStudy (theeric):

\[\huge \sqrt[3]{\frac{-512}{343}}\] I modified what you typed :)

OpenStudy (anonymous):

\[\sqrt[3]{\frac{-512}{343}} =\sqrt[3]{\frac{-512}{343}} \] \[\sqrt[3]{\frac{(-8) \times (-8) \times (-8)}{(7)\times(7)\times(7)}} = \frac{-8}{7}\]

OpenStudy (anonymous):

Note that the following expression can be re-written as:\[\large \bf \sqrt[3]{\frac{ (-2^3)^3 }{ 7^3 }}=\frac{ \sqrt[3]{(-2^3)^3} }{ \sqrt[3]{7^3} }=??\]

OpenStudy (theeric):

It's tough at first, but very pretty, huh?

OpenStudy (theeric):

It looks like you have the solution before you, @vishwasdevnani . If you haven't memorized your cubes, you'd just want to guess, and then see how close you are. Like, if you guessed \(\sqrt[3]{(-9)^3}\) for the numerator, you see that you had \(\sqrt[3]{-729}\), which is greater than \(\sqrt[3]{-512}\) and so you'd have to guess a smaller number, like \(\sqrt[3]{(-8)^3}\), which is the numerator.

OpenStudy (cggurumanjunath):

-8/7